Encyclopedia Astronautica
2008.01.19 - ISS On-Orbit Status 01/19/08


Saturday -- off-duty day for CDR Whitson, FE-1 Malenchenko and FE-2 Tani except for housekeeping and voluntary work.

CDR Whitson & FE-2 Tani started out with the daily reading of SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight) experiment data accumulated during the night, for logging and filling in questionnaire entries in the SLEEP session file on the HRF-1 laptop for downlink. (To monitor the crewmembers' sleep/wake patterns and light exposure, Dan and Peggy wear a special Actiwatch device which measures the light levels encountered by him as well as his patterns of sleep and activity throughout the Expedition. The log entries are done within 15 minutes of final awakening for seven consecutive days, as part of the crew's discretionary 'job jar' task list.)

Peggy and Dan performed their part of the regular weekly station cleaning (Yuri having done his share yesterday). ("Uborka", usually done on Saturdays, includes removal of food waste products, cleaning of compartments with vacuum cleaner, damp cleaning of the Service Module (SM) dining table, other frequently touched surfaces and surfaces where trash is collected, as well as the CDR's sleep station with a standard cleaning solution; also, fan screens and grilles are cleaned to avoid temperature rises. Special cleaning is also done every 90 days on the HEPA (high-efficiency particulate air) bacteria filters in the Lab.)

Tani also performed his daily status check on the BCAT-3 (Binary Colloidal Alloy Test-3) science payload, running by itself in Node-2 since 12/13/07 (briefly interrupted for EVA-13 photo support). (The status check, conducted on the last image taken by the DCS 760 digital still camera which is controlled by EarthKAM software on an A31p laptop, is to verify proper image focus and camera alignment. The SSC (Station Support Computer) is taking photography of the phase separation occurring in the BCAT Sample 3, with the photo flash going off every half hour.)

At ~9:05am EST, the crewmembers held their regular weekly planning conference (WPC) with the ground, discussing next week's "Look-Ahead Plan" (prepared jointly by MCC-H and TsUP/Moscow timeline planners) via S-band/audio, reviewing the monthly calendar, upcoming activities, and any concerns about future on-orbit events.

Later, Yuri completed today's routine maintenance of the SOZh system (Environment Control & Life Support System, ECLSS) in the Service Module (SM), including ASU toilet facilities systems/replaceables.

The crewmembers performed their regular 2.5-hr physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (CDR, FE-2), TVIS treadmill (FE-1), RED (CDR, FE-2) and VELO bike with bungee cord load trainer (FE-1).

Afterwards, Peggy Whitson copied the exercise data file to the MEC (Medical Equipment Computer) laptop for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

Yuri again had some time for transferring cargo delivered on Progress M-62/27P to the ISS and stowing it.

Conjunction Advisory: MCC-Houston ballistics specialists are monitoring a conjunction with orbital debris (Object #28099, an old Russian Strela ICBM body, NATO designation: SS-11 Sego), with TCA (time of closest approach) tomorrow morning at ~10:36am EST. Current predictions (Update #8) show a miss distance of ~20.6 km, which falls within the Notification Box according to Flight Rule (B4-101). Russian Ballistics has been notified. Prime DAM TIG (Debris Avoidance Maneuver Time of Ignition), based on current prediction, will be tomorrow morning at 8:18am EST for a TCA of 10:36am.

Weekly Science Update (Expedition Sixteen -- Week 13)

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): ALTEINO instrument has been rotated by the crew on 1/8 and memory card has been exchanged. Photos of the new configuration have been taken and placed on OCA for down-link. On 1/9 data has been downlinked and unexpected file size has been reported (partial recording). Analysis from the pictures showed that the grounding strap of the ALTEINO instrument got loose, leading to incorrect behavior of the instrument. After TsUP-Moscow request, the grounding cable has been fastened again and ALTEINO re-activated on 1/15.

ANITA: Completed.

BCAT-3 (Binary Colloidal Alloy Test 3): 'Here is our re-cap from last Monday's (1/14) ops to photo all samples then run sample 4. Seeing all the samples was very interesting and spurred the planned 1/24 sample 5 ops. We thank those involved in keeping EarthKAM going for BCAT!! Samples 1, 2, 3 and 6 all follow the same phase-separation sequence (as in the sample 3 movie he sent along), yet samples 4 and 5 are different and hence even more interesting. A scientific publication describing these findings is being prepared.'

CARDIOCOG-2: Completed.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): 'Dan, we are waiting to get the new launch date so we can schedule your second in-flight session to meet the R-21 to R-14 requirement. This activity will be hard scheduled when we get a firm launch date.'

CFE (Capillary Flow Experiment): Reserve.

CSI-2/CGBA (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus): In progress.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): In progress.

EarthKAM (Earth Knowledge Acquired by Middle School Students): Complete.

ELITE-S2 (Elaboratore Immagini Televisive - Space 2): Planned.

EPO (Educational Payload Operations): Reserve.

ETD (Eye Tracking Device): In progress.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): 'InSPACE-2 completed its nominal science this week and initiated an additional set of runs to further investigate pulsed-field structures at a lower pulse frequency. The fourth and highest concentration sample was successfully tested to conclude the nominal science. In total 17 runs have been completed. Our early observation of the video of the nominal science indicates that the chain structures aggregate laterally into a formation of thicker chains and voids, with a characteristic separation wavelength which depends on field strength and volume fraction. The chain aggregates are columns spanning the length of the cell, as opposed to ellipsoidal structures observed on the ground. The additional science runs provided the first glimpse of ellipsoidal structures with sharp pointed ends. A very interesting pulse behavior of the structures was observed. The aggregate structures pulsed lengthwise with the magnetic field and pulsed laterally in a wavy pattern was fascinating to watch.'

Integrated Immune: In progress.

KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Complete.

MISSE (Materials ISS Experiment): Ongoing.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 'Pirs'.

MULTIGEN-1: MULTIGEN-1 samples will be downloaded on STS-122 (1E).

MSG-SAME (Microgravity Science Glovebox): Complete.

NOA-2 (Nitric Oxide Analyzer): Planned.

NUTRITION/REPOSITORY: ' Dan, thank you for your voluntary participation in a stand-alone Repository session. Your attention in keeping to the time constraints was much appreciated!'

PMDIS (Perceptual Motor Deficits in Space): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): 'Peggy & Dan, we have added an Actiwatch download/initialization activity to the task list. The Actiwatches will stop taking data on 1/26 if not performed before then.'

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): In progress.

Swab (Characterization of Microorganisms & Allergens in Spacecraft): In progress.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.

CEO (Crew Earth Observation): Through 1/15 the ground has received a total of 10,857 frames of ISS/CEO imagery for review and cataloging. 'We have identified and are evaluating excellent imagery of our requests for: South Georgia Island; Gulf of Fonseca; and Perth, Australia (some of the best we've ever seen). We have also noted with delight your excellent imagery of cities at night. PAO is using one of your Dallas-Fort Worth, TX views. Your dramatic, oblique view of the Great Sand Dunes National Park and Preserve, CO will be published on NASA/GSFC's Earth Observatory website this weekend. Your view beautifully captures the context and extent of this spectacular dune field banked against the Sangre de Cristo Mountains. What a nice catch! Thanks for your fine work.'

CEO photo targets uplinked for today were Patagonian Glaciers, S. America (ISS orbit track brought the crew over the central Patagonian mountains. Looking to the left of track as the station passed over the mountains for glaciers and icefields near the summit ridgelines), and Jarvis Island, equatorial Pacific (weather was predicted to be clear over Jarvis Island, which is located roughly halfway between the Hawaiian and Cook Islands. Jarvis is an uninhabited coral island, but during the 19th century it was a mining location for bird guano. The island is now a U.S. National Wildlife Refuge (for seabirds, shorebirds, and marine fauna). Looking slightly to the left of track for the island.

More... - Chronology...

Home - Browse - Contact
© / Conditions for Use